1.   A human tissue cell has a diameter of the order 1.0µm. Assuming each cell occupies a cube of 1.0 µm on one side, estimate the number of cells in 1.0cm3 of tissue.


1
Expert's answer
2021-05-21T10:05:24-0400

Solution.

"d=1.0\mu m=10^{-6}m;"

"V=1.0cm^3=10^{-6}m^3;"

"V_1=d^3;"

"V_1=(10^{-6})^3=10^{-18}m^3;"

"N=\dfrac{V}{V_1};"

"N=\dfrac{10^{-6}}{10^{-18}}=10^{12};"

Answer:"N=10^{12}."